3. A point charge Qa = 3uC is located at point (0, 0) and another point Qb = - 7uC has the coordinates of (2, 3).



a) Find the electric force on charge Qa.



b) Is the force attractive or repulsive? Why?

1
Expert's answer
2022-05-19T16:40:17-0400

(a)

"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9\frac{3*10^{-6}*7*10^{-6}}{(2^2+3^2)}=1.45*{10^{-2}}\: \rm N"

(b) attraction.